# Explainable AI for computational pathology identifies model limitations and tissue biomarkers

### v2

@@ Abstract
Introduction: Deep learning models hold great promise for digital pathology, but their opaque decision-making processes undermine trust and hinder clinical adoption. Explainable AI methods are essential to enhance model transparency and reliability.
Methods: We developed HIPPO, an explainable AI framework that systematically modifies tissue regions in whole slide images to generate image counterfactuals, enabling quantitative hypothesis testing, bias detection, and model evaluation beyond traditional performance metrics. HIPPO was applied to a variety of clinically important tasks, including breast metastasis detection in axillary lymph nodes, prognostication in breast cancer and melanoma, and IDH mutation classification in gliomas. In computational experiments, HIPPO was compared against traditional metrics and attention-based approaches to assess its ability to identify key tissue elements driving model predictions.
Results: In metastasis detection, HIPPO uncovered critical model limitations that were undetectable by standard performance metrics or attention-based methods. For prognostic prediction, HIPPO outperformed attention by providing more nuanced insights into tissue elements influencing outcomes. In a proof-of-concept study, HIPPO facilitated hypothesis generation for identifying melanoma patients who may benefit from immunotherapy. In IDH mutation classification, HIPPO more robustly identified the pathology regions responsible for false negatives compared to attention, suggesting its potential to outperform attention in explaining model decisions.
Conclusions: HIPPO expands the explainable AI toolkit for computational pathology by enabling deeper insights into model behavior. This framework supports the trustworthy development, deployment, and regulation of weakly-supervised models in clinical and research settings, promoting their broader adoption in digital pathology.
@@ Introduction
Digital pathology has emerged as a transformative force in medicine, ushering in an era where computational methods can augment and enhance the diagnostic and prognostic capabilities of pathologists. By digitizing whole slide images (WSIs) of tissue specimens, this field has opened new avenues for applying advanced machine learning techniques to analyze complex histological patterns and features. The potential impact of computational pathology is far-reaching, promising to improve diagnostic accuracy, standardize interpretation, and uncover novel biomarkers that may inform personalized treatment strategies [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15].
@@
Building on these advances, the development of pathology foundation models represents a leap forward in computational pathology. These models have significantly expanded the field’s capabilities by providing powerful, generalizable tools for analyzing histological features [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. They are trained on vast collections of histology image patches in a self-supervised manner, enabling them to learn general-purpose representations of histological features. By focusing on patches, foundation models can efficiently process the large size and complexity of WSIs while capturing key local patterns. The representations learned by these models can be utilized in downstream tasks by incorporating patch-level embeddings as inputs for downstream models. For specimen-level prediction tasks, these embeddings are often integrated into an attention-based multiple instance learning (ABMIL[28]) framework. In this approach, the foundation model provides patch-level feature representations, and ABMIL aggregates and prioritizes relevant patches to generate specimen-level predictions. This combination has quickly become the standard for leveraging foundation models in specimen-level classification, offering a scalable and effective way to utilize their pre-trained knowledge for complex clinical and research applications.
Unlike traditional image classification methods, ABMIL models rely on specimen-level labels, eliminating the need for exhaustive pixel-level annotations [29]. This flexibility has enabled ABMIL to excel across a range of tasks, including cancer detection [30, 31], diagnosis [32, 33, 34, 35], identification of primary cancer origin [36], grading [31, 37, 38], genomic aberration detection [39, 40, 41, 42], molecular phenotyping [43, 44, 45], treatment response prediction [46, 47, 48], and prognostication [49, 50, 47, 51]. The integration of foundation models with ABMIL has proven particularly effective, combining the generalization capabilities of pre-trained models with the task-specific focus of ABMIL to advance computational pathology.
However, the widespread adoption of ABMIL models in clinical settings is hindered by challenges in model interpretability and trustworthiness [52, 53, 9, 10, 54, 55, 56]. Explainable AI methods have been proposed to address these challenges by providing insights into model behavior and identifying features that may influence predictions. Techniques such as LIME [57], SHAP [58], and SmoothGrad [59] offer strategies to estimate feature contributions in input images. However, these methods often rely on linear or additive approximations, which can oversimplify complex model behaviors and fail to align with actual model predictions [60, 61].
Attention mechanisms have emerged as a widely used approach for interpreting ABMIL models, as evidenced by their extensive use in recent studies [28, 46, 36, 50, 62, 63, 22, 64, 45, 65, 66, 67, 68, 69]. Although attention highlights regions of interest within WSIs, it does not directly quantify the influence of these regions on model predictions [70, 71]. This gap between attention and predictive influence can lead to misinterpretation, undermining trust in model decisions and limiting clinical applicability [72, 73, 74, 75]. Addressing this limitation requires methods that explicitly evaluate the contribution of specific regions to model outputs.
One such approach was introduced by Zeiler and Fergus, occluding parts of an image with gray boxes and observing how the model’s predictions change [76]. This approach can help identify which parts of an image are most relevant to a model’s decision. However, when applied to pathology, this method has limitations. Covering regions of WSIs with gray patches can cause erratic behavior in the model, as these unnatural patterns are not encountered during training, leading to interpretations that may be difficult to trust. While identifying regions to occlude in natural images often aligns with human intuition, this becomes more challenging in WSIs, where multiple regions of interest may collectively influence predictions. Additionally, occlusion methods only indicate the importance of regions but do not address whether the identified features are sufficient to drive the model’s predictions.
To better understand how specific tissue features influence model predictions, generating realistic counterfactual WSIs is essential. Counterfactual WSIs allow us to explore “what if” scenarios by systematically altering tissue features and observing the impact on model behavior. For example, resizing a tumor region could help determine the model’s sensitivity to tumor size, while modifying the number and distribution of tumor-infiltrating lymphocytes might reveal their role in influencing predictions, such as for prognosis or treatment response.
However, creating reliable counterfactual WSIs presents significant challenges. First, precise segmentation of tissue structures is necessary to enable targeted alterations without affecting unrelated regions. Second, robust inpainting techniques are required to fill in altered regions realistically, ensuring the resulting WSI appears biologically plausible. Although recent advances in generative AI show promise for creating synthetic images [77], applying these methods at the gigapixel scale for WSIs remains a technical challenge. The ability to generate realistic and valid counterfactual WSIs for clinical use has yet to be proven.
To address these challenges, we introduce HIPPO (Histopathology Interventions of Patches for Predictive Outcomes), an explainable AI method designed to enhance trust in ABMIL models and provide deeper insights into their decision-making processes. HIPPO goes beyond traditional attention-based interpretations by quantitatively assessing the impact of specific tissue regions on model predictions. HIPPO also leverages properties of multiple instance learning models to generate valid WSIs through the occlusion or inclusion of individual or groups of patches, simulating targeted interventions to understand how different histological features influence ABMIL model outputs.
@@
We validate HIPPO across three distinct tasks: metastasis detection, prognostication, and IDH mutation classification. In the context of metastasis detection, we evaluate five foundation models in pathology using the CAMELYON16 dataset [30]. This dataset includes expert-annotated labels and thus allows us to rigorously assess HIPPO’s ability to align with established ground truth and demonstrate its practical utility. Our analysis uncovers model-specific limitations and biases that would have remained hidden using attention mechanism alone. We reveal that some models rely heavily on extratumoral tissue for metastasis detection, while others are surprisingly insensitive to small tumor regions. These findings highlight the importance of rigorous model evaluation beyond standard performance metrics and underscore the potential of HIPPO in identifying when and why models might fail. Building on this validation, we then aplied HIPPO to cancer prognostication in breast cancer and cutaneous melanoma datasets from The Cancer Genome Atlas. Our results demonstrate that HIPPO can identify tissue regions more strongly associated with prognosis compared to those highlighted by attention. Strikingly, we find that high-attention regions can sometimes have counterintuitive effects on prognostic predictions, further emphasizing the limitations of relying solely on attention for model interpretation. By quantitatively assessing the impact of tumor-infiltrating lymphocytes (TILs) on model predictions, we confirm that the models have captured the known prognostic significance of TILs in both breast cancer and melanoma. This ability to link model behavior to established biological knowledge is crucial for building trust in AI-driven prognostic tools. Finally, we applied HIPPO to IDH mutation classification in glioma and found that HIPPO outperformed attention in identifying the histological regions driving false negatives. By providing more precise and interpretable insights into model predictions, HIPPO enabled a clearer understanding of the reasons behind misclassifications.
Beyond model interpretation, HIPPO opens new possibilities for virtual hypothesis generation in computational pathology. We showcase this potential by using HIPPO to simulate the effects of autologous TIL therapy in melanoma patients. By digitally replicating TILs in specimens with poor prognosis, we demonstrate a proof-of-principle approach for identifying patients who potentially might benefit from this immunotherapy. This application illustrates how HIPPO can bridge the gap between computational predictions and clinically actionable insights.
@@
As computational pathology continues to advance, the need for robust, interpretable, and trustworthy AI models becomes increasingly critical. HIPPO represents a significant step forward in this direction, offering a powerful tool for uncovering the strengths, limitations, and potential biases of ABMIL models in pathology. By providing a more comprehensive understanding of model behavior, HIPPO not only enhances the interpretability of existing models but also paves the way for developing more reliable and clinically relevant AI tools in pathology. As we demonstrate across multiple applications, from metastasis detection to prognostic modeling and mutation classification, HIPPO has the potential to accelerate the translation of computational pathology into clinical practice, ultimately improving patient care and outcomes.
@@ Results
HIPPO is a specimen-level perturbation toolkit that explains weakly-supervised models in computational pathology (Fig. 1a). The fundamental goal of HIPPO is to explore counterfactual (i.e., “what if”) scenarios that are infeasible to realize in actual tissue samples. For instance, it would be impractical to directly manipulate the tumor microenvironment of a tissue specimen to understand its effect on a prognostic model. Instead, we can digitally modify a WSI that simulates this intervention. HIPPO enables virtual interventions through the occlusion or inclusion of single or multiple patches, utilizing the resulting ABMIL model predictions as counterfactual outcomes. HIPPO provides quantitative insights into how specific tissue alterations might impact pathological assessments through the lens of the AI model. These assessments can include but are not limited to, patient prognosis, treatment response prediction, metastasis detection, inference of spatial transcriptomics, gene mutation detection, and microsatellite instability identification. Applying HIPPO to ABMIL models enables researchers, regulators, and clinicians to elucidate model behavior and to better understand the relationship between patch-level tissue characteristics and model outputs.
Traditional approaches to digital interventions in medical imaging often require precise segmentation of objects for occlusion or inclusion [76, 78], as well as sophisticated inpainting techniques to maintain image integrity [79, 80, 81, 82]. Alternatively, generative AI can generate counterfactual images [83, 84], but the quality of the generated images has not been thoroughly evaluated for histopathology. These manual or AI-assisted methods can introduce covariate shifts when imperfectly executed [85], potentially leading to unreliable model predictions. The key insight for HIPPO is based on how data flows through ABMIL models. A WSI is treated as a bag of permutation-invariant patches, where the number and order of patches are allowed to vary [28]. Thus, an intervention can be achieved through two primary perturbation mechanisms: (1) removing specific patches, effectively excising tissue from the input specimen, or (2) including specific patches, simulating the addition of new tissue into the specimen. HIPPO leverages unique properties of multiple instance learning models to facilitate the generation of counterfactual images bypassing the complexities of direct image manipulation by creating hypothetical scenarios such as the introduction or removal of tumor patches or regions of tumor-infiltrating lymphocytes (TILs) from a patient’s specimen. Understanding when ABMIL models alter their predictions due to interventions provides quantitative insights into their decision making process, revealing important features and potential biases learned.
HIPPO offers hypothesis-driven and data-driven methods for intervention selection:
@@
HIPPO-knowledge: choosing a region based on prior knowledge or a well-defined hypothesis and quantifying its effect by removing it from specimens or adding it to specimens without that region and measuring the change in model outputs (Fig. 1a).
HIPPO-attention: quantifying the effect of high attention regions by removing the high attention regions and measuring the change in model outputs (Fig. 1b).
@@
HIPPO-search-high-effect: a greedy search algorithm to identify the regions that maximally drive a prediction. This can be used to identify regions necessary for a model’s output (Fig. 1c).
HIPPO-search-low-effect: a greedy search algorithm to identify the regions that lead to the smallest change in model predictions. This can be used to identify the regions that do not change model predictions and, therefore, are not necessary for the output of a model. After removing unnecessary regions, the remaining regions can be considered sufficient for the model prediction (Fig. 1c).
Hypothesis-driven interventions leverage prior knowledge of relevant patches, which can be identified with the assistance of pathologists or annotation tools. In cases where prior information is unavailable or when important features are sub-visual (e.g., in prognostic prediction tasks), data-driven interventions can automatically identify important patches de novo. Unlike traditional attribution methods that rely on additive approximations, HIPPO supports multi-patch interventions, enabling the exploration of how sets of patches collectively drive non-linear effects on predictions.
It is important to note that HIPPO generates counterfactual predictions based on the capabilities of the underlying AI model. Consequently, if the AI model has limitations in capturing complex biological relationships, these limitations will inherently influence HIPPO’s explanations. Furthermore, the resolution of HIPPO’s interventions is constrained to patches, which may restrict its ability to capture finer-grained details. Despite these constraints, HIPPO offers a novel and powerful approach for interrogating the biological factors driving predictions, shedding light on the decision-making process of AI models in unprecedented ways.
@@
With the advent of digital pathology foundation models, it is important to evaluate model robustness, generalizability, and potential biases and understand their limitations. Here, we showcase the breakthroughs made possible by HIPPO in rigorously evaluating models built on top of foundation models for breast metastasis detection and prognosis prediction tasks. We compare five foundation models in metastasis detection and identify model-specific limitations and biases. We also use HIPPO to study the effects of tissue components on prognostic models, demonstrating how HIPPO’s capabilities surpass attention in identifying low and high-risk drivers. We also measure the effect of TILs on breast cancer and melanoma patient prognosis and demonstrate digitally that autologous TILs improve predicted prognosis in a subset of melanoma patients, marking exciting progress in the field.
@@
Metastasis detection is a well-studied task, with well-defined features (i.e., tumor cells) that drive the label of whether or not a specimen contains metastasis. In a clinical setting, it is critical that metastases are identified; a false negative is unacceptable. Recent studies have shown that ABMIL models have strong performance in metastasis detection[20]. However, previous studies have also found that computer vision models can make the correct predictions for the wrong reasons, such as short-cut features or spurious correlations [86, 87]. Thus, the degree to which AI models rely on the tumor regions remains to be seen, even for a relatively straightforward task like tumor detection. Understanding this is critical to elucidate the strengths and limitations of ABMIL models for metastasis detection, including potential biases.
@@
To evaluate this, we trained several ABMIL models for breast metastasis detection using the CAMELYON16 dataset [30] (Fig. 2a). Several pathology foundation models have recently emerged, demonstrating near-human levels in metastasis detection. Here we consider five pathology foundation models (UNI [20], REMEDIS [16], Phikon [18], CTransPath [17], and RetCCL [19]). We trained five ABMIL models for each foundation model to distinguish whether or not a specimen contained metastasis. Similar to previously reported results [20], UNI achieved a mean balanced accuracy of 0.982, REMEDIS 0.922, Phikon 0.907, CTransPath 0.858, and RetCCL 0.745. (Fig. 2b, Supplementary Table 1). For HIPPO explainability experiments, we used the best-performing model (out of 5 random initializations) on the test set for each foundation model. The best UNI model achieved balanced accuracy of 1.00, REMEDIS 0.949, Phikon 0.955, CTransPath 0.885, and RetCCL 0.769 (Supplementary Table 2).
@@
In this dataset, expert pathologists have produced high resolution annotations of metastatic tumor regions. This allows us to use the HIPPO-knowledge method to determine whether metastatic regions are necessary for detecting breast cancer metastasis. Specifically, for patients who were positive for metastasis, we created counterfactual examples by removing all patches that intersected with the tumor annotations. If we assume that a model must rely on tumor-containing patches to make a prediction, an accurate model would always return a negative prediction when all tumor containing patches have been removed. In the following, we use specificity to quantify model behavior. Specificity is defined as the ratio of true negatives to all negative samples. In this set of counterfactuals, all specimens were negative, so the specificity represented the proportion of correct negative predictions by the models.
@@
We compared model predictions before and after removing patches containing tumor. Notably, the UNI-based model exhibited the lowest specificity (0.73) in these counterfactual examples despite achieving the highest balanced accuracy on the original test set (1.00). This discrepancy was particularly pronounced in counterfactual specimens that originally contained macrometastases (specificity 0.59), suggesting that the UNI-based ABMIL model uses tissue outside of the tumor region to drive positive metastasis predictions. The REMEDIS-based model exhibited a similar trend, with a specificity of 0.77 in counterfactuals derived from macrometastases. In contrast, the other models showed less dependence on extratumoral tissue (sensitivity of Phikon-based, 0.86; CTransPath-based, 0.92; RetCCL-based, 0.88), indicating that their predictions are primarily driven by tumor epithelial cells rather than other tissue components (Fig. 2c).

### v3

@@ Abstract
Deep learning models show promise in digital pathology, but their opaque decision-making processes undermine trust and limit their widespread clinical adoption. To address this challenge, we present HIPPO, an explainable AI method for analyzing weakly-supervised multiple-instance learning (MIL) models that are widely used in whole slide image analysis. HIPPO constructs counterfactual whole slide images by systematically removing or adding selected tissue regions, providing a principled way to quantify how specific histologic areas influence model predictions under the MIL framework. This capability enables rigorous model interpretation, quantitative hypothesis testing, bias detection, and performance evaluation that extend beyond standard metrics. We demonstrate HIPPO’s capabilities across numerous models performing clinically important tasks, such as breast metastasis detection in axillary lymph nodes, prognostication in breast cancer and melanoma, and IDH mutation classification in gliomas. In metastasis detection, HIPPO uncovered critical model limitations that were undetectable by standard performance metrics or attention-based methods. For prognostic prediction, HIPPO outperformed attention by providing more nuanced insights into tissue elements influencing outcomes. In a proof-of-concept study, HIPPO facilitated hypothesis generation for identifying melanoma patients who may benefit from immunotherapy. In IDH mutation classification, HIPPO more robustly identified the pathology regions responsible for false negatives compared to attention, suggesting its potential to outperform attention in explaining model decisions. In summary, HIPPO expands the explainable AI toolkit for computational pathology by enabling deeper insights into model behavior. This framework supports the trustworthy development, deployment, and regulation of weakly-supervised models in clinical and research settings, promoting their broader adoption in digital pathology.
@@ Introduction
Digital pathology has emerged as a transformative force in medicine, ushering in an era where computational methods can augment and enhance the diagnostic and prognostic capabilities of pathologists. By digitizing whole slide images (WSIs) of tissue specimens, this field has opened new avenues for applying advanced machine learning techniques to analyze complex histological patterns and features. The potential impact of computational pathology is far-reaching, promising to improve diagnostic accuracy, standardize interpretation, and uncover novel biomarkers that may inform personalized treatment strategies [van2021deep, echle2021deep, topol2019high, song2023artificialcpath, niazi2019digital, rakha2021current, cui2021artificial, morales2021artificial, tizhoosh2018artificial, tran2021deep, cifci2023ai, shmatko2022artificial, perez2024guide, unger2024systematic, bera2019artificial].
@@
Building on these advances, the development of pathology foundation models represents a leap forward in computational pathology. These models provide powerful, generalizable tools for analyzing histological features [azizi2022remedis, wang2022ctranspath, filiot2023phikon, wang2023retccl, chen2024uni, lu2024visualconch, xu2024whole, wang2024pathology, vorontsov2024virchow, zimmermann2024virchow2, filiot2024phikonv2, nechaev2024hibou, aben2024kaiko, vorontsov2024virchow, zimmermann2024virchow2]. Trained on vast collections of histology image patches in a self-supervised manner, they learn general-purpose representations of histological features. By focusing on patches, foundation models can efficiently process the large size of WSIs while capturing key local patterns. The resulting patch-level representations can then be used as inputs to downstream task-specific models. For specimen-level prediction tasks, these embeddings are often integrated into a multiple-instance learning (MIL) framework, such as attention-based MIL (ABMIL) architecture [ilse2018abmil]. In this setup, the foundation model provides patch-level embeddings, and ABMIL identifies relevant patches and assigns attention weights to aggregate them, finally obtaining specimen-level predictions (Fig. 1a). This combination has quickly become the standard for leveraging foundation models in specimen-level classification, offering a scalable and effective way to utilize their pre-trained knowledge for complex clinical and research applications. Unlike traditional image classifiers, MIL models only require labels for a bag of image patches (i.e., specimen-level labels) for training, enabling scalable application across diverse tasks [gadermayr2024multiple]. This flexibility has enabled MIL-based workflows to excel across a range of tasks, including cancer detection [bejnordi2017camelyon16, bulten2022panda], diagnosis [qiu2021attention, butke2021end, del2022constrained, del2021attention], identification of primary cancer origin [lu2021ai], grading [bulten2022panda, su2022attention2majority, li2021prostategrading], genomic aberration detection [saldanha2023self, dernbach2024dissecting, zheng2024predicting, schirris2021deepsmile], molecular phenotyping [valieris2024weakly, boehm2024oncotype, el2024regression], treatment response prediction [lipkova2022crane, jiang2023biology, mallya2024benchmarking], and prognostication [chen2021multimodal, chen2022porpoise, jiang2023biology, ammeling2023attention].
However, the widespread adoption of ABMIL models in clinical settings is hindered by challenges in model interpretability and trustworthiness [plass2023explainability, castro2020causality, tizhoosh2018artificial, tran2021deep, foote2021now, ghaffari2022adversarial, dawood2024buyer]. General-purpose explainable AI methods, such as LIME [ribeiro2016lime], SHAP [lundberg2017shap], and SmoothGrad [selvaraju2017grad], have been proposed to address these challenges by providing insights into model behavior and identifying features that may influence predictions. These methods often rely on linear or additive approximations, which can oversimplify complex model behaviors [lipton2018mythos, bilodeau2024impossibility]. In ABMIL models, attention scores, which are used as weights for aggregating patch embeddings, have emerged as a widely used approach for interpreting ABMIL models, as evidenced by their extensive use in recent studies [ilse2018abmil, lipkova2022crane, lu2021ai, chen2022porpoise, lu2021clam, xiong2023diagnose, wang2024pathology, niehues2023generalizable, el2024regression, graziani2022attention, cai2024pathologist, yao2024application, zhang2023attention, sehring2023leveraging]. Although attention highlights regions of interest within WSIs, it does not directly quantify the influence of these regions on model predictions [javed2022additivemil, nan2024establishing, jain2019attention]. This gap between attention and predictive influence can lead to misinterpretation, undermining trust in model decisions and limiting clinical applicability [liu2024attention, raff2024reproducibility, lin2023interventional, pocevivciute2020survey]. There is a pressing need for methods that explicitly evaluate how specific tissue regions contribute to model outputs.
One principled approach is to create counterfactual WSIs in which the subregion containing the feature of interest is ablated. Such counterfactual inputs would allow us to explore “what if” scenarios by observing the differential impact on model behavior. For example, simulating WSIs with larger or smaller tumor regions could help determine the model’s sensitivity to tumor size, while modifying regions with tumor-infiltrating lymphocytes might reveal their role in influencing predictions, such as for prognosis or treatment response. However, obtaining reliable counterfactual WSIs presents significant technical challenges. Occlusion-based methods (e.g., masking regions with gray boxes [zeiler2014visualizing]) can produce unnatural artifacts and cause erratic behavior in the model, as these patterns were not encountered during training, leading to interpretations that may be difficult to trust. To remove specific information from an image while keeping it realistic, generative models offer a potential solution [pozzi2024synthetic], but applying robust inpainting or image synthesis at gigapixel scales for WSIs to fill in the modified regions realistically remains a technical challenge. The ability to generate realistic and valid counterfactual WSIs for clinical use has yet to be proven. Consequently, practical and trustworthy counterfactual explanations for computational histopathology models remain a major unmet need.
To address these challenges, we introduce HIPPO (Histopathology Interventions of Patches for Predictive Outcomes), an explainable AI method designed to enhance trust in pathology models and provide deeper insights into their decision-making processes (Methods). HIPPO simulates counterfactual WSIs by modifying the set of patches that compose a slide, selectively removing regions or adding new ones, to probe how model predictions change under controlled perturbations (Fig. 1b). HIPPO leverages the permutation-invariant nature of MIL models, which treat WSIs as unordered bags of patches, allowing these counterfactual patch sets to serve as valid inputs without requiring synthetic image generation or inpainting. Since MIL models are permutation-invariant, these resampled patch collections remain valid inputs to models. By enabling targeted interventions on patch sets, HIPPO allows us to systematically analyze how different histological features influence MIL model outputs. Unlike attention mechanisms, which highlight correlated regions but cannot test causal relationships, HIPPO supports hypothesis-driven analyses by explicitly measuring whether a region is necessary, sufficient, or irrelevant for a prediction. This capability elevates interpretability from descriptive observation to empirical hypothesis testing. Compared to traditional attention-based interpretations, HIPPO provides direct, quantitative assessments of the impact of specific tissue regions on model predictions.
@@
We demonstrate how HIPPO enables our systematic understanding of the workings of a variety of pathology models performing clinical tasks: metastasis detection, prognostication, and IDH mutation classification. First, in the context of metastasis detection, we evaluate six foundation models in pathology using the CAMELYON16 dataset [bejnordi2017camelyon16]. This dataset includes expert-annotated tumor segmentations made by pathologists, providing ground-truth labels that allow us to rigorously determine whether regions deemed influential by the model correspond to established histopathologic ground truth. Throughout our metastasis detection experiments, we extensively use these expert-annotated tumor regions to validate HIPPO’s causal attributions, testing both the necessity and sufficiency of tumor regions for model predictions against this ground truth. Our analysis uncovers model-specific limitations and biases that would have remained hidden if relying solely on attention-based interpretations. We reveal that some models rely heavily on extratumoral tissue for metastasis detection, while others are surprisingly insensitive to small tumor regions. These findings highlight the importance of rigorous model evaluation beyond standard performance metrics and underscore the potential of HIPPO in identifying when and why models might fail. Second, building on this validation, we then applied HIPPO to cancer prognostication in breast cancer and melanoma datasets from The Cancer Genome Atlas (TCGA). Our results demonstrate that HIPPO can identify tissue regions more strongly associated with prognosis compared to those highlighted by attention. Strikingly, we find that high-attention regions can sometimes have counterintuitive effects on prognostic predictions, further emphasizing the limitations of relying solely on attention for model interpretation. By quantitatively assessing the impact of tumor-infiltrating lymphocytes (TILs) on model predictions, we confirm that the models have captured the known prognostic significance of TILs in both breast cancer and melanoma. This ability to link model behavior to established biological knowledge is crucial for building trust in AI-driven prognostic tools. Beyond model interpretation, HIPPO opens new possibilities for virtual hypothesis generation in computational pathology. Finally, we applied HIPPO to IDH mutation classification in glioma and found that HIPPO outperformed attention in identifying the histological regions driving false negatives. By providing more precise and interpretable insights into model predictions, HIPPO enabled a clearer understanding of the reasons behind misclassifications.
@@
As computational pathology continues to advance, the need for robust, interpretable, and trustworthy AI models becomes increasingly critical. HIPPO addresses this need by providing a systematic framework for uncovering the strengths, limitations, and potential biases of models. By offering a more complete characterization of model behavior, HIPPO enhances the interpretability of existing approaches and facilitates the development of more reliable and clinically meaningful pathology models. As demonstrated across multiple applications, including metastasis detection, prognostic modeling, and mutation classification, HIPPO can help accelerate the translation of computational pathology methods into clinical practice and, ultimately, support improved patient care.
@@ Results
HIPPO is a specimen-level perturbation framework that explains permutation-invariant MIL models in computational pathology (Fig. 1a). The primary goal of HIPPO is to explore counterfactual scenarios (e.g., “how would the model’s output change if the tumor microenvironment within a WSI were altered?”). Without relying on synthetic image generation, HIPPO performs such interventions through the exclusion or inclusion of single or multiple patches, leveraging the permutation-invariant structure of MIL models (Fig. 1b). The resulting changes in model predictions can then be interpreted as counterfactual outcomes. Through this process, HIPPO provides quantitative insights into how specific tissue subregions might impact pathological assessments of the AI model. These assessments can include but are not limited to patient prognosis, treatment response prediction, metastasis detection, inference of spatial transcriptomics, gene mutation detection, and microsatellite instability identification. Applying HIPPO to pathology models enables researchers, regulators, and clinicians to elucidate model behavior and to better understand the relationship between patch-level tissue characteristics and model outputs.
The key insight behind HIPPO lies in how data flows through MIL models. In these models, a WSI is treated as a bag of permutation-invariant patches, where the number and order of patches are allowed to vary. This formulation enables interventions through either (1) removing specific patches, effectively excising tissue from the input specimen, or (2) including specific patches, simulating the addition of new tissue into the specimen (Fig. 1b and Supplementary Fig. LABEL:fig:hippo-schematic-insert-delete). Since MIL models are essentially set functions, these modified patch collections remain valid model inputs and do not introduce spatial artifacts that could confuse models and confound interpretations (Supplementary Fig. LABEL:fig:supp_permutation_invariance) . This intervention approach enables us to explore hypothetical scenarios such as the exclusion or inclusion of tumor patches from a patient’s specimen. Understanding how the predictions of MIL models change due to interventions provides quantitative insights into their decision making process, revealing important features and potential biases learned.
Within the HIPPO framework, users can flexibly select which patches to include or exclude, guided either by prior hypotheses or by data-driven strategies. Accordingly, we categorize HIPPO usage into the following modes (see Methods and LABEL:sec:supplementary_methods for details):
@@
HIPPO-knowledge: choosing patches to add or remove based on prior knowledge or a well-defined hypothesis and quantifying its effect by removing it from specimens or adding it to specimens without that region and measuring the change in model outputs. (Fig. 1c). In other words, this hypothesis-driven approach assesses whether the areas considered to be important by experts truly affect model outputs.
HIPPO-attention: choosing a subregion to ablate based on the attention scores assigned by MIL models and quantifying its effect by removing it from specimens and measuring the change in model outputs (Fig. 1d). In other words, this method assesses whether the areas with attention scores are indeed influential for the model’s predictions.
HIPPO-search-positive-effect: a greedy search algorithm to identify the regions that most strongly drive a prediction to the positive direction. We apply this method to samples predicted as positive by the model to identify regions influential for their predictions (Fig. 1e). This data-driven intervention can automatically identify important patches from scratch without relying on prior hypotheses.
@@
HIPPO-search-negative-effect: a greedy search algorithm to identify the regions that most strongly drive a prediction to the negative direction. We apply this method to samples predicted as negative by the model to identify regions influential for their predictions.
@@
With the advent of digital pathology foundation models, it is important to evaluate model robustness, generalizability, and potential biases and understand their limitations. Here, we showcase the advance enabled by HIPPO in rigorously evaluating models built on top of foundation models for breast metastasis detection, prognosis prediction tasks, and IDH mutation detection task. We compare six foundation models in metastasis detection and identify model-specific limitations and biases. We also use HIPPO to study the effects of tissue components on prognostic models, demonstrating how HIPPO’s capabilities surpass attention in identifying low and high-risk drivers. We also measure the effect of TILs on breast cancer and melanoma patient prognosis and demonstrate digitally that autologous TILs improve predicted prognosis in a subset of melanoma patients, marking exciting progress in the field.
@@
Metastasis detection is a well-studied task with well-defined features (i.e., tumor cells) that drive the label of whether or not a specimen contains metastasis. In a clinical setting, it is critical that metastases are identified — a false negative is unacceptable. Recent studies have shown that MIL models have strong performance in metastasis detection[chen2024uni]. However, previous studies have also found that computer vision models can make the correct predictions for the wrong reasons, such as short-cut features or spurious correlations [geirhos2020shortcut, degrave2021shortcuts]. Thus, the degree to which AI models rely on the tumor regions remains to be seen, even for a relatively straightforward task like tumor detection. Understanding this is critical to elucidate the strengths and limitations of MIL models for metastasis detection, including potential biases.
@@
To evaluate this, we trained several ABMIL models for breast metastasis detection using the CAMELYON16 dataset [bejnordi2017camelyon16] (Fig. 2a and Methods). While HIPPO is applicable to any permutation-invariant MIL models, we focus on ABMIL models because they are the de facto standard in computational pathology among MIL approaches. Several pathology foundation models have recently emerged, demonstrating near-human levels in metastasis detection. Here we consider six pathology foundation models (UNI [chen2024uni], REMEDIS [azizi2022remedis], Phikon [filiot2023phikon], CTransPath [wang2022ctranspath], RetCCL [wang2023retccl], and Virchow2[vorontsov2024virchow, zimmermann2024virchow2]) . We trained five ABMIL models for each foundation model to distinguish whether or not a specimen contained metastasis. Similar to previously reported results [chen2024uni], UNI achieved a mean balanced accuracy of 0.982, REMEDIS 0.922, Phikon 0.907, CTransPath 0.858, RetCCL 0.745, and Virchow2 0.934 (Fig. 2b, Supplementary Table LABEL:table:supp_cam16_performance). For HIPPO explainability experiments, we used the best-performing model (out of 5 random initializations) on the test set for each foundation model. The best UNI model achieved balanced accuracy of 1.00, REMEDIS 0.949, Phikon 0.955, CTransPath 0.885, RetCCL 0.769, and Virchow2 0.959 (Supplementary Table LABEL:table:supp_cam16_performance_single_models).
@@
In this dataset, expert pathologists have produced high-resolution annotations of metastatic tumor regions. These expert-annotated ground-truth labels allow us to use the HIPPO-knowledge method to determine whether metastatic regions are necessary for detecting breast cancer metastasis, with the expert annotations serving as the ground-truth reference. Specifically, for patients who were positive for metastasis, we created counterfactual examples by removing all patches that intersected with the tumor annotations. If we assume that a model must rely on tumor-containing patches to make a prediction, an accurate model would always return a negative prediction when all tumor containing patches have been removed. In the following, we use specificity to quantify model behavior. Specificity is defined as the ratio of true negatives to all negative samples. In this set of counterfactuals, all specimens were negative, so the specificity represented the proportion of correct negative predictions by the models.
@@
We compared model predictions before and after removing patches containing tumor (Fig. 2c). Notably, the UNI-based model exhibited the lowest specificity (0.73) in these counterfactual examples despite achieving the highest balanced accuracy on the original test set (1.00). This discrepancy was particularly pronounced in counterfactual specimens that originally contained macrometastases (specificity 0.59), suggesting that the UNI-based ABMIL model uses tissue outside of the tumor region to drive positive metastasis predictions. The REMEDIS-based model exhibited a similar trend, with a specificity of 0.77 in counterfactuals derived from macrometastases. In contrast, the other models showed less dependence on extratumoral tissue (sensitivity of Phikon-based, 0.86; CTransPath-based, 0.92; RetCCL-based, 0.88; Virchow2-based, 0.92 ), indicating that their predictions are primarily driven by tumor epithelial cells rather than other tissue components.